OpenStudy (anonymous):

A bacteria culture begins with 13 bacteria which triple in amount every hour. How many bacteria exist in the culture at the 11th hour?

OpenStudy (tkhunny):

How many are there after one hour?

OpenStudy (anonymous):

I don't know. The answer choices are 13,312 26,624 767,637 2,302,911

OpenStudy (tkhunny):

I do not believe that. Read the problem statement again and tell me how many there are after one hour. Start with 13 and triple it.

OpenStudy (anonymous):

13 x 3 = 39 (hr 1) 39 x 3 = 117 (hr 2) 117 x 3 = 351 (hr 3) 351 x 3 = 1053 (hr 4) 1053 x 3 = 3159 (hr 5) 3159 x 3 = 9477 (hr 6) 9477 x 3 = 28431 (hr 7) 28431 x 3 = 85293 (hr 8) 85293 x 3 = 255879 (hr 9) 255879 x 3 = 767637 (hr 10) 767637 x 3 = 2,302,911 (hr 11)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Sorry I don't have the technical know how to write that in a usable formula

OpenStudy (tkhunny):

t is the number of hours since we started. \(f(t) = 13\cdot 3^{t}\) Find f(11).

OpenStudy (anonymous):

That's wrong for future references

OpenStudy (tkhunny):

Actually, that's exactly correct, in case you need an immediate reference. Perhaps, sometime in the future, arithmetic will work differently. For now, you probably should go with how it works, today.